# Stereo Hand-Object Reconstruction for Human-to-Robot Handover

### v1

@@ III-D Handover with stereo hand-object reconstruction
We integrate StereoHO into a pipeline for human-to-robot handovers. We use two cameras on either side of the robot to maximise the visible area of the hand and object.
To perform stereo hand-object reconstruction, we first obtain the necessary inputs from the RGB image. For each view, a hand-object detector estimates the bounding boxes of both the hand and the object. The object bounding box is used as a prompt for object segmentation, whereas we use the bounding box of the hand to crop the image around the hand. The cropped image is used as input to a hand capture module that predicts the shape of the hand and the pose of the wrist. We remove the background from the cropped image by using the object segmentation mask and the segmentation mask of the hand, extracted as a silhouette from the shape of the hand.
We use the cropped image, the pose of the wrist, and the segmentation masks to reconstruct the pointclouds of the hand and the object using StereoHO. We ensure multi-view consistency by reprojecting the predicted hand and object pointcloud $\mathcal{P}_{H}$ and $\mathcal{P}_{O}$ back to the RGB images and keep only the points that are within the segmentation masks $M_{H}$ and $M_{O}$, respectively. Since the reconstruction can change at each timestep, it can significantly affect the grasp estimation quality. Therefore, we only update the predicted pointclouds if there is an improvement in reconstruction quality. We compute the Intersection over Union (IoU) of the convex hull of the pointclouds with the segmentation masks, respectively, and only accept the updated reconstruction if the IoU is greater than the previous best IoU.
III-E Object receiving and delivery
Given the pointclouds predicted by our hand-object reconstruction model, we estimate a set of 6-DoF grasps, $\mathcal{G}=\{g_{n}\in SE(3)\}_{n=1}^{N}$, on the object pointcloud $\mathcal{P}^{\prime}_{O}$. The predicted grasps $\mathcal{G}$ are in the hand wrist coordinate frame and is transformed using the predicted wrist pose $T_{HB}$ to the robot base coordinate frame. To ensure the safety of the human, we filter out any grasps that are colliding with the hand point cloud $\mathcal{P}^{\prime}_{H}$
@@ III-E Object receiving and delivery
The robot control selects the grasp $g$ closest to the current gripper 6D pose at every timestep. We define a standoff position 15 cm away from the grasp point in the direction of the grasp. Once the robot reaches the standoff position, the robot enters a close-loop grasping stage and closes its fingers when the gripper reaches the selected grasp.
After the grasp $g$ is executed, the robot retracts for 15 cm back to the standoff position. The robot then moves to a predefined delivery location and opens the gripper fingers to place the object on the table. Finally, the robot returns to its home position ready for the next handover.
III-F Implementation details
For hand-object reconstruction, we set $w_{c}=0.25$ for the index $c$ corresponding to empty space and $w_{c}=0.75$ for the rest of the codebook as the majority of the space surrounding the hand and object is empty. We use ResNet-18 [29] pre-trained on ImageNet[30] as our image encoder. During the handover, we use hand-object detector [31] to obtain hand-object bounding boxes, Fast Segment Anything [32] for object segmentation, FrankMocap [33] for hand capture, and 6-DoF GraspNet [14] for grasp estimation.
@@ IV-A Datasets
We train our hand-object reconstruction model on the ObMan [11] dataset and test on the DexYCB [17] dataset. ObMan is a fully simulated hand-object dataset using the MANO hand model [34] with ShapeNet [35] objects. The object classes include bottles, bowls, cans, jars, knifes, cellphones, cameras and remote controls. DexYCB is a hand-object dataset captured using an 8 camera setup in a lab environment. 20 objects from the YCB-Video dataset [36] were used including cans, boxes and various everyday objects.
@@ IV-B Handover setup
We follow an existing benchmarking protocol [9] that includes a setup of the robotic and perception systems, and 288 handover configurations to execute. The setup includes a 6-DoF Universal Robots UR5 robotic arm, equipped with a Robotiq 2F-85 2-finger gripper for grasping the object, and two Intel Realsense D435 cameras mounted on tripods on either side of the robotic arm pointing at the workspace of the robot which is a wooden table with the target delivery location marked on the surface. The protocol configurations are executed by 4 participants who hand over 3 disposable cups and 1 plastic wine glass, either empty or filled with rice, using 3 different grasp types and at 3 handover locations (72 configurations for each participant). To assess generalisation to different novel objects, each participant execute another 15 configurations for 5 objects at the 3 different handover locations using 1 grasp type (a total of 60 configurations). These additional objects vary in size, topology, and appearance (e.g., reflectiveness), and include a tea box, a cracker box, a tape, a screwdriver, and a CD (see Fig. 5).
Figure 5: Objects used for the handover experiments: three cups and a drinking glass, empty or filled with rice, from an existing benchmarking protocol [9] (left image), and additional household objects (right image), such as tea box (O1), cracker box (O2), tape (O3), screwdriver (O4), and CD (O5), to assess generalisation.
We compare our human-to-robot handover method with the stereo RGB method [12] designed specifically for container-like objects [^2] Approval for this research was obtained from the Queen Mary Ethics of Research Committee Ref:QMERC20.525 on 19/10/2021. Consent from each human subject was obtained prior to the experiments.. For [12], the objects start on the table at the centre of the delivery target and participants are asked to pick up the object and give it to the robot. The object on the table is first detected using Mask R-CNN [37] pretrained on the COCO dataset [38]. The shape is then estimated using LoDE [16]. The mask is tracked with SiamMask [39] in subsequent frames and the 3D object centroid is triangulated from the 2D centroid in the left and right view. The container is assumed to be upright at all times throughout the handover process and hand keypoints are estimated using OpenPose [40] per frame. Grasp points are generated along the height of the container and grasp points that are close to the hand keypoints in height are removed. The robot starts moving once the object is lifted off the table. For our method, since we assume the human is already holding the object, we ask the participants to hold the object behind a starting line indicated by tape on the wooden table. The robot starts moving towards the object once the human moves the object beyond the starting line.
Figure 6: Comparison of single view and stereo hand object reconstructions on DexYCB [17]. For the single-view setting, each reconstruction corresponds to the image on the same row. For the stereo setting, the same reconstruction is shown from two viewpoints. Our method yields less noisy reconstruction by introducing segmentation masks as input. In the stereo setting, our method improves both the hand and object reconstructions by combining the predictions from individual views.
@@ IV-C Performance measures
The less time required, the more efficient the method is, and the higher the efficiency score $\gamma$. The score is scaled based on the expected minimum time required $\eta=1,000$ ms and the maximum time required $\tau=15,000$ ms.
@@
IV-D Hand-object reconstruction from RGB
Since our reconstruction method combines individual single-view predictions from the left and right views to form the stereo prediction, we evaluate our method in both single-view and stereo settings.
We compare StereoHO with IHOI [19] in the single-view setting and with SVHO [10] in the stereo setting. IHOI is an SDF-based method that predicts the shape of the object conditioned on the pose of the hand, and is trained end-to-end. SVHO is a sparse multi-view method that reconstructs hand-object shape and can be easily adapted to use only two views. For a fair comparison, we train our method, IHOI, and SVHO, only on the synthetic data of ObMan, and evaluate the methods using DexYCB [17]. We select the two views closest to the stereo views of our setup and compare the reconstruction quality (see Fig. 6). For the single-view setting, we aggregate results from both the left and right views. We separate the objects in the testing set into seen and unseen categories based on the similarity in 3D shape between the training and testing objects.
@@ IV-E Human-to-robot handovers
TABLE I: Comparing handover performance between [12] which performs grasping with fixed rotation and our method which performs 6-DoF grasping with rotation. We validate on (Left) container-like objects from the CORSMAL benchmark with and without rice, and (Right) additional household objects. Our method is able to handover container-like objects and additional household objects, while [12] cannot handle the additional household objects as the perception module does not recognize non-container objects and hence cannot perform the handover.
Score Ref. Container-like objects Grasp types Filling Household objects
All C1 C2 C3 C4 TOP MID BOT E FR All O1 O2 O3 O4 O5
Delivery location $\delta$ [12] 0.45 0.52 0.58 0.33 0.38 0.47 0.35 0.53 0.40 0.51
Ours 0.48 0.26 0.65 0.59 0.42 0.45 0.45 0.55 0.47 0.49 0.18 0.18 0.11 0.06 0.25 0.31
Efficiency $\gamma$ [12] 0.29 0.31 0.35 0.26 0.25 0.31 0.29 0.29 0.79 0.78
Ours 0.10 0.12 0.09 0.10 0.09 0.10 0.10 0.10 0.10 0.10 0.10 0.09 0.07 0.07 0.14 0.13
Delivered mass $\mu$ [12] 0.78 0.77 0.81 0.77 0.78 0.81 0.81 0.84 0.79 0.78
Ours 0.74 0.53 0.87 0.84 0.73 0.66 0.69 0.88 0.76 0.72
Grasping success $G$ [12] 0.79 0.77 0.81 0.79 0.80 0.82 0.72 0.84 0.79 0.80
Ours 0.75 0.55 0.88 0.84 0.73 0.66 0.69 0.89 0.76 0.74 0.70 0.50 0.66 0.75 0.75 0.83
Delivery success $D$ [12] 0.67 0.73 0.77 0.50 0.68 0.72 0.53 0.76 0.61 0.73
Ours 0.66 0.36 0.86 0.84 0.59 0.63 0.58 0.78 0.69 0.63 0.58 0.41 0.58 0.50 0.58 0.83
Best scores are highlighted in bold for each column and each performance measure. KEY – TOP: top, MID: middle, BOT: bottom, E: empty, FR: full rice, Ref.: reference. C1: small white cup, C2: red cup, C3: beer cup, C4: wine glass, O1: tea box, O2: cracker box, O3: tape, O4: screwdriver, O5: CD.
We first evaluate the methods with the objects from the CORSMAL benchmark (Tab. I). Both [12] and our method achieve over 75% grasping success and 65% delivery success. This shows that assuming a cylindrical prior and a mostly upright pose throughout the task is reasonable when handing over cups and drinking glasses [12]. However, [12] performs grasping with a fixed rotation, forcing the subjects to adapt to the robot’s gripper pose (Fig. 8). On the contrary, the 6-DoF grasping of our method allows the subject to be more natural, but at the cost of being computationally more expensive and achieving a lower efficiency score (0.10).
Figure 8: Our method allows the robot to perform grasping with 6-DoF, resulting in more natural handovers, compared to [12] that performs grasping with fixed rotations.
Results per object show that our method can successfully handover transparent objects with minimal degradation in performance. The grasping and delivery scores for beer cup (transparent object) only differ by 0.04 and 0.02, respectively, compared to those of red cup (opaque object). Our method also performs better on larger objects and detects the object already handheld. Smaller objects are hard to detect for our method due to occlusions from the hand and background clutter from the clothes of the participants. On the contrary, [12] performs better on smaller objects because the detection starts when the object is unoccluded and on the table with a uniform background. Results by grasp type show that our method performs best when objects are handed over with the bottom grasp. Note that this grasp type is more common in the training set of our reconstruction model. The middle grasp leaves the least amount of space on the container for the robot to grasp, decreasing the performance of [12].
@@
Tab. I also shows that our method achieved 70% grasping success and 58% delivery success when handing the additional household objects over to the robot, whereas [12] cannot detect them as they are not present in the COCO dataset [38]. Specifically, our method can successfully receive objects with challenging shape, such as thin objects (the screwdriver), hollow objects (the tape), and reflective objects (the CD). Failures occur when the reconstruction of the object shape is inaccurate. For example, the robot often mistakenly try to grasp the cracker box from the wider side, which is not possible for the width of the gripper (see Fig. 9). Another mode of failure occurs when the robot loses track of the hand due to occlusions by the object or the robot arm.
Figure 9: Example handovers. Each column shows a single configuration from initialisation to object delivery.
@@ V Conclusion
In this paper, we tackled the task of joint hand-object reconstruction using wide baseline stereo RGB cameras. We showed that our reconstruction method outperforms existing methods in single-view and stereo settings. Moreover, we showed that our method enables a robot to receive empty and filled containers, and can generalize to other household objects. Future work includes improving the speed of handovers and the quality of hand-object reconstructions.

### v2

@@ III-D Handover with stereo hand-object reconstruction
We integrate StereoHO into a pipeline for human-to-robot handovers (Fig. 4) and assume that i) the handover task starts with the person already holding the object and ii) two cameras are placed on either side of the robot, maximising the visible area of the hand and object.
For each frame of the approaching phase, StereoHO reconstructs the pointclouds of the hand and the object using image crops, pose of the wrist, and segmentation masks from each view. Therefore, for each camera view and for each frame, a hand-object detector [32] estimates the bounding boxes of both the hand and the object, $b_{H},b_{O}\in\mathbb{R}^{4}$. We use Fast Segment Anything [33] to segment the object ($M_{O}$) within the object bounding box [33] and we crop the image around the hand using $b_{H}$. From the image crop, we use FrankMocap [34] to predict the pose of the wrist, $T_{H}$ and the segmentation mask of the hand, $M_{H}$, after extracting the silhouette from the hand shape. We use both segmentation masks to remove the background from the image crop. We triangulate the centroid of the object mask and verify the reprojection error in each view is less than 5 pixels. Note that the pipeline does not perform the handover (failure case) if, within 10 s, no hand-object pair is detected in either of the views or the triangulation is not validated. Since the reconstruction quality can vary across frames, significantly affecting the estimated robot grasp, we compute the Intersection over Union (IoU) of the convex hull of the pointclouds with the segmentation masks, respectively. We update the pointclouds if there is an improvement in reconstruction quality (i.e., IoU $>$ IoU).
@@
To enable the robot to receive the object, we use 6-DoF GraspNet [15] to estimate a set of $N$ candidate grasps, $\mathcal{G}=\{g_{n}\in SE(3)\}_{n=1}^{N}$, on the object pointcloud $\mathcal{P}^{\prime}_{O}$. To ensure human safety, we construct a 3D bounding box of the size of the gripper around each grasp, and we filter out any grasps with points from the hand point cloud $\mathcal{P}^{\prime}_{H}$ in its bounding box. Using the predicted wrist pose in the robot coordinate frame, $T_{HB}=T_{WB}T_{HW}$ where $T_{WB}$ is the robot base transform, we transform the predicted grasps from the hand-wrist to the robot coordinate frame. For each timestep, the robot control selects the grasp $g$ closest to the current gripper 6-DoF pose. We define a standoff position 15 cm away from the grasp point in the direction of the grasp. Once the robot reaches the standoff position, the robot enters a close-loop grasping stage and closes its fingers when the gripper reaches the selected grasp. After grasping, the robot retracts to the standoff position, delivers the object at a predefined location on the table, and returns to its initial position.
III-E Parameter settings
We set the T-SDF dimension to $D=128$. The codebook size is $C=512$ with embedding size $S=128$. For $\mathcal{L}_{ae}$, we set $\beta=1.0$. For $\mathcal{L}_{ce}$, we set $w_{c}=0.25$ for the index $c$ corresponding to empty space and $w_{c}=0.75$ for the rest of the codebook. The robot base transform $T_{WB}$ is obtained using the hand-eye calibration process [35]. We estimate $N=200$ grasp candidates for each handover.
@@ IV-A Datasets
We train StereoHO on the ObMan [13] dataset and test it on DexYCB [18]. The training set of ObMan contains 87,190 images of synthetic handheld objects rendered on top of real images. Objects were selected from ShapeNet [36] based on 8 categories (Fig. 5), such as bottles, bowls, cans, jars, knifes, cellphones, cameras and remote controls, and grasped by the MANO hand model [37]. DexYCB is a multi-view dataset with videos of a person manipulating one of 20 everyday objects selected from YCB-Video [38].
@@ IV-B Handover setup
We follow a benchmarking protocol [4] that includes a setup of the robotic and perception systems, and 288 handover configurations to execute. The setup includes a 6-DoF Universal Robots UR5 robotic arm, equipped with a Robotiq 2F-85 2-finger gripper for grasping the object, and two Intel Realsense D435i cameras mounted on tripods on either side of the robotic arm pointing at the workspace of the robot which is a table with the delivery target marked. The protocol configurations are executed by 4 participants who hand over 3 cups and 1 plastic wine glass, either empty or filled with rice, using 3 different grasp types at 3 handover locations (72 configurations for each participant) [^3] Approval for this research was obtained from the Queen Mary Ethics of Research Committee Ref:QMERC20.525 on 19/10/2021. Consent from each human subject was obtained prior to the experiments.. The protocol instructs to initially place the object on the table at the centre of the delivery target, and then a person picks up and hands the object over to the robot. On the contrary, we assume the object to be already held by the person for our pipeline and therefore we asked participants to hold the object behind a starting line indicated by tape on the wooden table. The robot starts moving towards the object once the human moves the object beyond the starting line. To assess generalisation to different novel objects outside of the benchmark, each participant execute another 24 configurations for 8 household objects at the 3 different handover locations using 1 grasp type (a total of 96 configurations). These additional objects vary in size, topology, and appearance (Fig. 5).
Figure 5: Synthetic object types used for training StereoHO, and objects for testing hand-object reconstruction and handover experiments. For handovers: 3 cups and 1 glass, empty or filled with rice, from a benchmarking protocol [4] (left), and 8 household objects (right) to assess generalisation.
Figure 6: Comparing single-view and stereo hand object reconstructions on DexYCB [18]. For single-view, each reconstruction corresponds to the image on the same row. For stereo, the same reconstruction is shown from two viewpoints. Our method yields less noisy reconstructions and improves both the hand and object shape by combining the single view predictions.
@@ IV-C Performance measures
The less time required, the more efficient the method is, and the higher the efficiency score $\gamma$. The score is scaled based on the expected minimum time required $\eta=1$ s and the maximum time allowed $\tau=15$ s.
@@
IV-D Hand-object reconstruction from RGB
Since our reconstruction method combines two individual single-view predictions to form the stereo prediction, we evaluate our method in both single-view and stereo settings. We compare StereoHO with IHOI [20] in the single-view setting and with SVHO [27] in the stereo setting. IHOI is an SDF-based method that predicts the shape of the object conditioned on the pose of the hand, and is trained end-to-end. SVHO is a sparse multi-view method that reconstructs hand-object shape and can be easily adapted to use only two views. For a fair comparison, we train our method, IHOI, and SVHO, only on the synthetic data of ObMan, and evaluate the methods using DexYCB [18]. We select the two views closest to the stereo views of our setup and compare the reconstruction quality (Fig. 6). For the single-view setting, we aggregate results from both the left and right views. We separate the testing set into seen and unseen categories based on the similarity in 3D shape between the training and testing objects (Fig. 5).
@@ IV-D Hand-object reconstruction from RGB
TABLE I: Comparison of performance results between hand-object reconstruction methods integrated in a system for human-to-robot handovers of unknown and diverse household objects with varying physical properties (shape, mass, transparency).
Score Method DoF RGB D DC SC HH CORSMAL containers Household objects
Avg C1 C2 C3 C4 Avg O1 O2 O3 O4 O5 O6 O7 O8 Avg C3 C4
$\delta$ CB [6] 3 0.45 0.52 0.58 0.33 0.38 - - - - - - - - - 0.31 0.05 0.57
DB 6 - - - - - 0.06 0.05 0.00 0.05 0.06 0.11 0.00 0.00 0.20 0.01 0.03 0.00
ClearGrasp [11] 6 - - - - - 0.01 0.00 0.00 0.00 0.00 0.00 0.05 0.00 0.08 0.00 0.00 0.00
StereoHO 6 0.48 0.26 0.65 0.59 0.42 0.18 0.18 0.11 0.06 0.25 0.31 0.09 0.18 0.23 0.55 0.65 0.45
$\gamma$ CB [6] 3 0.29 0.31 0.35 0.26 0.25 - - - - - - - - - 0.20 0.16 0.24
DB 6 - - - - - 0.29 0.34 0.27 0.30 0.27 0.37 0.27 0.10 0.38 0.13 0.10 0.16
ClearGrasp [11] 6 - - - - - 0.14 0.15 0.05 0.06 0.14 0.15 0.28 0.14 0.16 0.09 0.10 0.08
StereoHO 6 0.10 0.12 0.09 0.10 0.09 0.21 0.09 0.07 0.07 0.14 0.13 0.39 0.44 0.35 0.09 0.09 0.09
$\mu$ CB [6] 3 0.78 0.77 0.81 0.77 0.78 - - - - - - - - - - - -
StereoHO 6 0.74 0.53 0.87 0.84 0.73 - - - - - - - - - - - -
$G$ CB [6] 3 0.79 0.77 0.81 0.79 0.80 - - - - - - - - - 0.83 0.66 1.00
DB 6 - - - - - 0.41 0.16 0.16 0.50 0.50 0.58 0.50 0.08 0.83 0.08 0.08 0.08
ClearGrasp [11] 6 - - - - - 0.15 0.00 0.00 0.08 0.25 0.16 0.41 0.00 0.33 0.20 0.33 0.08
StereoHO 6 0.75 0.55 0.88 0.84 0.73 0.75 0.50 0.66 0.75 0.75 0.83 0.75 1.00 0.75 0.91 1.00 0.83
$D$ CB [6] 3 0.67 0.73 0.77 0.50 0.68 - - - - - - - - - 0.50 0.08 0.91
DB 6 - - - - - 0.20 0.08 0.16 0.33 0.25 0.25 0.00 0.00 0.58 0.04 0.08 0.00
ClearGrasp [11] 6 - - - - - 0.05 0.00 0.00 0.00 0.00 0.00 0.16 0.00 0.25 0.00 0.00 0.00
StereoHO 6 0.66 0.36 0.86 0.84 0.59 0.50 0.41 0.58 0.50 0.58 0.83 0.08 0.58 0.41 0.83 1.00 0.66
Results of our proposed method highlighted in gray. Objects highlighted based on their transparency (transparent, opaque, mixed). Best scores highlighted in bold for each column and each performance measure. KEY – DoF: degrees of freedom, RGB: colour input, D: depth input, DC: depth completion, SC: stereo camera, HH: handheld object assumption, Avg: average, C1: small white cup, C2: red cup, C3: beer cup, C4: wine glass, O1: tea box, O2: cracker box, O3: tape, O4: screwdriver, O5: CD, O6: cloth, O7: cap, O8: spray bottle; $\delta$: delivery location, $\gamma$: efficiency, $\mu$: delivered mass, $G$: grasping success, $D$: delivery success, : not used, : used; -: not applicable.
@@ IV-E Human-to-robot handovers
We compare StereoHO with a depth-based baseline (DB), ClearGrasp [11], and the CORSMAL baseline (CB) [6]. DB uses the pointcloud observed from the left camera. ClearGrasp is a depth completion method to handle transparent objects using depth-based sensing. For fairness, DB and ClearGrasp are integrated in the same handover pipeline. CB [6] is a stereo-based pipeline specifically designed for handovers of rotationally symmetric container objects. Unlike our pipeline, CB assumes the object is initially placed on the table, following the original protocol [4], and is upright throughout the entire handover process. The object is detected using Mask R-CNN [39], pretrained on COCO [40], and the object shape is estimated with a multi-view iterative fitting algorithm given a cylindrical prior [17]. For each frame, hand keypoints are estimated using OpenPose [41] and the 3D object centroid is triangulated from the 2D centroid in the left and right views after tracking the object mask in each view with SiamMask [42]. Grasp points are generated along the height of the container and those ones close to the hand keypoints on the vertical axis are removed. The robot starts moving once the object is lifted off the table.
Figure 8: Comparing hand (red) and object (green) reconstructions, overlaid on the pointcloud obtained by the camera.
Tab. I compares the handover performance of the methods. For the container objects from the CORSMAL benchmark [4], both CB and our method achieve over 75% grasping success and 65% delivery success. This shows that assuming a cylindrical prior and a mostly upright pose throughout the task is reasonable when handing over cups and drinking glasses [6]. When handing over general household objects, our method achieved 75% grasping success on the opaque objects and 91% grasping success on the transparent containers, significantly outperforming DB and ClearGrasp. Comparing the hand and object shape reconstructed by each method (Fig. 8), DB is easily affected by transparency of the object and noisy depth measurements, leaving holes in the reconstructed shape. ClearGrasp can fill in the missing depth value but the estimated value is often inaccurate and noisy. Our method produces a complete shape which helps the grasp estimation method to produce successful grasp candidates for the handover.
Figure 9: Examples of successful human-to-robot handovers of household objects using StereoHO, and failure cases due to grasping failure or spilled filling.
@@
Overall, we showed that our method can perform human-to-robot handovers on general household objects (Fig. 9), outperforming depth-based methods and performing on par on containers with a method designed specifically for the handover of containers. Specifically, our method can successfully receive objects with challenging shapes, such as thin (screwdriver), deformable (cloth), and reflective objects (CD), whereas previous stereo-based method fail to generalise to non-container objects, and depth-based methods fail to receive transparent objects. Failures occur when the reconstruction of the object shape is inaccurate. For example, the robot often mistakenly try to grasp the cracker box from the wider side (Fig. 9), which is larger than the gripper width. Failure can also occur when the robot loses track of the hand due to occlusions by the object or the robot arm itself.
@@ V Conclusion
In this paper, we proposed a joint hand-object reconstruction method to enable human-to-robot handovers using wide baseline stereo RGB cameras. Our reconstruction method outperforms existing methods in both single-view and stereo settings. Handover experiments showed that our method enables a robot to receive empty and filled containers, and can generalize to other household objects. Future work includes improving the speed of handovers and the quality of hand-object reconstructions.